OpenStudy (anonymous):

Situation: A 40 gram sample of a substance thats used for drug research has a k-value of 0.1476. N=Noe^-kt No=initial mass( at time t=0) N=mass at time t k=a positive constant that depends on the substance itself and on the units used to measure time t=time, in days Find the substance's half-life,in days.Round your answer to the nearest tenth.

OpenStudy (anonymous):

these questions are all the same. what you are looking for is the time when N=No/2. if you sub that in you get No/2=N0e^-kt. the No's cancel and solving for t you get t=-ln(0.5)/k

OpenStudy (anonymous):

as you can see the answer is independent of the initial mass which is a characteristic of half-lives

OpenStudy (anonymous):

i have to find the substances half like in days

OpenStudy (anonymous):

so as you can see the t value when calculated will give you the time in days as the k value adjusts for this so if you plug the solution into a calculator you get 4.69611... rounded to the nearest tenth you get 4.7 days

OpenStudy (anonymous):

formula for half life=(ln2)/k =(ln2)/0.1476=4.7 approximately

OpenStudy (anonymous):

thankssssss

OpenStudy (anonymous):

can u guys help me with the next question?

OpenStudy (anonymous):

4.7